# Taming Hyperparameter Sensitivity in Data Attribution: Practical Selection Without Costly Retraining

### v1

@@ 1 Introduction
Data attribution methods quantify the contribution of individual training samples to a machine learning model’s predictions or overall performance [21]. These methods have been successfully applied to a range of data-centric tasks in modern AI, including training data selection [36], mislabel detection [21], and copyright compensation [7]. Recently, this field has witnessed a surge of novel methods [10, 17, 29, 11, 28, 6, 3, 34], each promising improvements in efficiency, efficacy, or theoretical soundness.
@@
Hyperparameter tuning is particularly challenging in data attribution due to the high computation costs associated with common evaluation metrics. Standard evaluation metrics—such as Leave-One-Out (LOO) correlation [21] and Linear Datamodeling Score (LDS) [28]—require retraining models on subsets of the full training dataset for hundreds or even thousands of times. A conventional hyperparameter search under such evaluation metrics is prohibitively expensive in practice.
@@ 2 Background
In this section, we provide the necessary background to contextualize our empirical study of hyperparameter sensitivity in Section 3 and the subsequent case analysis in Section 4.
@@ 2.1 Data Attribution Methods and Their Hyperparameters
Consider a multi-class classification problem with dataset $S=\{(x_{i},y_{i})\}_{i=1}^{n}\subset\mathcal{Z}$, where $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, inputs $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$, and labels $y_{i}\in\mathcal{Y}$. We use a parameterized hypothesis function $h(\cdot,\theta):\mathcal{X}\rightarrow\Delta_{|\mathcal{Y}|-1}$ with parameters $\theta\in\mathbb{R}^{p}$, which typically outputs class probabilities via a softmax layer. The cross-entropy loss for a single example $z=(x,y)$ is denoted by $L(z,\theta)=\ell(h(x,\theta),y)$. The empirical risk at $\theta$ over a subset $A\subset S$ of size $a$ is $R_{A}(\theta):=\frac{1}{a}\sum_{z\in A}L(z,\theta)$. For a given test example $z^{\prime}=(x^{\prime},y^{\prime})$, following Park et al. [28], we define the model output function as $f(z^{\prime},\theta):=\ln\frac{p(z^{\prime},\theta)}{1-p(z^{\prime},\theta)}$, where $p(z^{\prime},\theta)=\exp(-L(z^{\prime},\theta))$ is the model’s predicted probability for the correct class $y^{\prime}$. We denote the parameter that minimizes $R_{S}$ as $\theta_{S}^{*}$.
@@
A data attribution method (or an attributor) is a function $\tau:\mathcal{Z}\times S\rightarrow\mathbb{R}$ that, for a given test example $z^{\prime}\in\mathcal{Z}$, assigns $\tau(z^{\prime},z_{i})$ to each training point $z_{i}\in S$. A popular attributor is the Influence Function (IF) attributor, introduced by Koh and Liang [21], which computes the influence of a training example $z_{i}\in S$ on a test example $z^{\prime}\in\mathcal{Z}$ as
$\tau_{\mathrm{IF}}(z^{\prime},z_{i}):=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}H_{S}^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*}),$ (1)
where $H_{S}:=\nabla_{\theta}^{2}R_{S}(\theta_{S}^{*})$ is the Hessian of the empirical risk $R_{S}$. To bypass the prohibitive computation of $H_{S}$, they also propose using efficient approaches to approximate $H_{S}^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*})$, including conjugate gradient (CG) [26] and stochastic estimation method LiSSA [1]. For non-convex models, $H_{S}$ is usually replaced with its regularized version $H_{S}+\lambda I_{p}$ in case of singularity [21], where $\lambda>0$ is a regularization term and $I_{p}$ is the identity matrix of size $p$. To further handle non-convexity and reduce computational costs, Grosse et al. [11] draw from second-order optimization literature [25] and introduce the Generalized Gauss Newton matrix (GGN) and empirical Fisher Information Matrix (FIM) to approximate $H_{S}$, which are guaranteed to be positive semi-definite (or even positive definite when a small regularization $\lambda I_{p}$ is included), and only involve first-order differentiation. We formally define the Influence Function attributor with FIM (IFFIM) as
$\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z_{i}):=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*}),$ (2)
@@
$F_{S}:=\frac{1}{n}\sum_{z_{i}\in S}[\nabla_{\theta}L(z_{i},\theta_{S}^{*})\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}]=\mathbb{E}_{z_{i}\sim S}[\nabla_{\theta}L(z_{i},\theta_{S}^{*})\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}].$ (3)
@@
$\tau_{\mathrm{IFFIM},\lambda,P}(z^{\prime},z_{i}):=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-1}P^{\top}\nabla_{\theta}L(z_{i},\theta_{S}^{*}).$ (4)
Another popular method, TracIn [29], computes the influence of $z_{i}$ on $z^{\prime}$ by tracing the entire optimization process based on stochastic gradient descent. Mathematically, it works as if substituting $H_{S}^{-1}$ with the identity matrix, while taking average across training model checkpoints weighted by the learning rate: $\tau_{\mathrm{TracIn}}(z^{\prime},z_{i}):=\sum_{t}\eta_{t}\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}\nabla_{\theta}L(z_{i},\theta_{t}),$ where $\theta_{t}$ and $\eta_{t}$ are the parameters and learning rate at checkpoint $t$.
@@ 2.2 Evaluation Metrics for Data Attribution
$c_{s}(\tau,z^{\prime},\overline{A}):=\mathrm{SpearmanCorr}\bigg{(}\{f(z^{\prime},\theta^{*}_{A_{j}})\}_{j\in[s]},\big{\{}\sum_{z\in A_{j}}\tau(z^{\prime},z)\big{\}}_{j\in[s]}\bigg{)},$ (5)
@@ 3.1 Experimental Setup
We evaluate a range of data attribution methods, including IF and its variants (CG and LiSSA) [21], TracIn [29], TRAK [28], and LoGra [6], introduced in Section 2.1.
@@
We perform empirical evaluations on three standard benchmark settings: MNIST [24] with a multilayer perceptron (MLP), CIFAR-2 with ResNet-9 [23, 13], and WikiText2 [27] with the GPT2 [30], which consists of both image and text data with various model sizes. We use the implementation from the dattri library [8] for the datasets and models.
@@
We consider both the common hyperparameters introduced in Section 2.1 and some critical method-specific hyperparameters. For TRAK, we experiment with regularization, projection-dimension, and training-epoch. For TracIn, we search for projection-dimension, normalization, and checkpoint-selection. For IF, we analyze regularization and training-epoch, as well as max-iteration for the CG variant, and scaling and recursion-depth for the LiSSA variant. For LoGra, we search for regularization, projection-dimension, and training-epoch. We design the search space for each hyperparameter around its default value proposed by the original papers. The detailed definitions and the search space of the hyperparameters are stated in Appendix A.1.
@@ 3.2 Experimental Results
((c)) CIFAR-2+ResNet9. Attributor: TRAK. HP: projection-dimension, regularization.
@@
((e)) CIFAR-2+ResNet9 & MNIST+ MLP. Attributor: TRAK. HP: training-epoch.
@@
  • Most attribution methods are sensitive to certain hyperparameters. Small changes in hyperparameters can lead to large variations in LDS performance, highlighting the sensitivity of most methods to hyperparameter choices. For example, Figure 1(a) shows that increasing regularization from 1e-4 to 1e-2 for IF (CG) leads to a 3$\times$ increase in LDS.
@@ 4 A Case Study on the Regularization Term in Influence Function
4.1 Analysis: Hyperparameter Selection of $\lambda$ in Influence Function
We study the problem of selecting $\lambda$ using the IFFIM version of influence function, $\tau_{\mathrm{IFFIM},\lambda}$, as defined in Eq. 2.
@@ 4.1 Analysis: Hyperparameter Selection of $\lambda$ in Influence Function
Recall Definition 2.1, where we have denoted LDS as $c_{s}(\tau,z^{\prime},\overline{A})$ for an attributor $\tau$, a test sample $z^{\prime}$, and the sampled subset $\overline{A}$. The hyperparameter selection problem can be formulated as the following optimization problem:
@@
$c_{p}(\tau,z^{\prime}):=\mathrm{PearsonCorr}\bigg{(}\{f(z^{\prime},\theta^{*}_{A})\}_{A\subset S,|A|=a},\big{\{}\sum_{z\in A}\tau(z^{\prime},z)\big{\}}_{A\subset S,|A|=a}\bigg{)}.$ (6)
Note that Spearman correlation is equivalent to the Pearson correlation of the rank-transformed variables. Therefore, both statistics measure monotonic relationships and are generally expected to be highly correlated [9]. The new objective, Population Pearson LDS, is differentiable and replaces the sampled subsets $\overline{A}$ with the expectation, which makes it easier to analyze. Specifically, we can define the partial derivative of $c_{p}$ with respect to $\lambda$ as
$\dot{c_{p}}(\lambda;z^{\prime}):=\frac{\partial c_{p}(\tau_{\mathrm{IFFIM},\lambda},z^{\prime})}{\partial\lambda}.$ (7)
@@
In our hyperparameter sensitivity experiments in Section 3, we observe that, empirically, the LDS with respect to the regularization term $\lambda$ often has a uni-modal and concave shape. This observation turns our attention from the original maximization problem to the analysis of the partial derivative $\dot{c_{p}}(\lambda;z^{\prime})$. We first show in the following Lemma 4.2 that our empirical observation about the shape of the LDS as a function of $\lambda$ is not a coincidence.
@@
Assuming the test sample $z^{\prime}$ follows a similar distribution as the training data, under some technical assumptions deferred to Section B.3, [^7] This involves non-trivial technical assumptions that do not always strictly hold. But we have made intuitive and empirical justifications about them. with high probability over the choice of test sample $z^{\prime}$, there exists some $C>0$ such that for $0<\lambda<C$,
@@
Lemma 4.2 suggests that for a range of small $\lambda$ starting from 00, $c_{p}$ as a function of $\lambda$ is increasing. With the observation regarding the shape of the LDS curve, ideally, we can solve for the optimal hyperparameter $\lambda^{*}$ as the stationary point with $\dot{c_{p}}(\lambda^{*};z^{\prime})=0$. However, the partial derivative $\dot{c_{p}}(\lambda^{*};z^{\prime})$ still involves quantities $\theta_{A}^{*}$ for $A\subset S$, which require model retraining.
A sufficient condition for positive derivative.
Instead of directly solving for the stationary point, we derive a sufficient condition (Theorem 4.3) for the derivative to be positive, which leads to a surrogate indicator that does not require model retraining (see Definition 4.6 in Section 4.2).

### v2

@@ 1 Introduction
Data attribution methods quantify the contribution of individual training samples to a machine learning model’s predictions or overall performance [22, 9]. These methods have been successfully applied to a range of data-centric tasks in modern AI, including training data selection [37], mislabel detection [22], and copyright compensation [7]. Recently, this field has witnessed a surge of novel methods [11, 18, 30, 12, 29, 6, 3, 35], each promising improvements in efficiency, efficacy, or theoretical soundness.
@@
Hyperparameter tuning is particularly challenging in data attribution due to the high computation costs associated with common evaluation metrics. Standard evaluation metrics—such as Leave‑One‑Out (LOO) correlation [22] and Linear Datamodeling Score (LDS) [29]—require retraining models on subsets of the full training dataset for hundreds or even thousands of times. A conventional hyperparameter search under such evaluation metrics is prohibitively expensive in practice.
@@ 2 Background
In this section, we provide the necessary background to contextualize our empirical study of hyperparameter sensitivity in Section˜3 and the subsequent case analysis in Section˜4.
@@ 2.1 Data Attribution Methods and Their Hyperparameters
Consider a multi-class classification problem with dataset $S=\{(x_{i},y_{i})\}_{i=1}^{n}\subset\mathcal{Z}$, where $\mathcal{Z}=\mathcal{X}\times\mathcal{Y}$, inputs $x_{i}\in\mathcal{X}\subset\mathbb{R}^{d}$, and labels $y_{i}\in\mathcal{Y}$. We use a parameterized hypothesis function $h(\cdot,\theta)\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\Delta_{|\mathcal{Y}|-1}$ with parameters $\theta\in\mathbb{R}^{p}$, which typically outputs class probabilities via a softmax layer. The cross-entropy loss for a single example $z=(x,y)$ is denoted by $L(z,\theta)=\ell(h(x,\theta),y)$. The empirical risk at $\theta$ over a subset $A\subset S$ of size $a$ is $R_{A}(\theta)\mathrel{\mathop{\ordinarycolon}}=\frac{1}{a}\sum_{z\in A}L(z,\theta)$. For a given test example $z^{\prime}=(x^{\prime},y^{\prime})$, following Park et al. [29], we define the model output function as $f(z^{\prime},\theta)\mathrel{\mathop{\ordinarycolon}}=\ln\frac{p(z^{\prime},\theta)}{1-p(z^{\prime},\theta)}$, where $p(z^{\prime},\theta)=\exp(-L(z^{\prime},\theta))$ is the model’s predicted probability for the correct class $y^{\prime}$. We denote the parameter that minimizes $R_{S}$ as $\theta_{S}^{*}$.
@@
A data attribution method (or an attributor) is a function $\tau\mathrel{\mathop{\ordinarycolon}}\mathcal{Z}\times S\rightarrow\mathbb{R}$ that, for a given test example $z^{\prime}\in\mathcal{Z}$, assigns $\tau(z^{\prime},z_{i})$ to each training point $z_{i}\in S$. A popular attributor is the Influence Function (IF) attributor, introduced by Koh and Liang [22], which computes the influence of a training example $z_{i}\in S$ on a test example $z^{\prime}\in\mathcal{Z}$ as
$\tau_{\mathrm{IF}}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}H_{S}^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*}),$ (1)
where $H_{S}\mathrel{\mathop{\ordinarycolon}}=\nabla_{\theta}^{2}R_{S}(\theta_{S}^{*})$ is the Hessian of the empirical risk $R_{S}$. To bypass the prohibitive computation of $H_{S}$, they also propose using efficient approaches to approximate $H_{S}^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*})$, including conjugate gradient (CG) [27] and stochastic estimation method LiSSA [1]. For non-convex models, $H_{S}$ is usually replaced with its regularized version $H_{S}+\lambda I_{p}$ in case of singularity [22], where $\lambda>0$ is a regularization term and $I_{p}$ is the identity matrix of size $p$. To further handle non-convexity and reduce computational costs, Grosse et al. [12] draw from second-order optimization literature [26] and introduce the Generalized Gauss Newton matrix (GGN) and empirical Fisher Information Matrix (FIM) to approximate $H_{S}$, which are guaranteed to be positive semi-definite (or even positive definite when a small regularization $\lambda I_{p}$ is included), and only involve first-order differentiation. We formally define the Influence Function attributor with FIM (IFFIM) as
$\tau_{\mathrm{IFFIM},\lambda}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}(F_{S}+\lambda I_{p})^{-1}\nabla_{\theta}L(z_{i},\theta_{S}^{*}),$ (2)
@@
$F_{S}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n}\sum_{z_{i}\in S}[\nabla_{\theta}L(z_{i},\theta_{S}^{*})\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}]=\mathbb{E}_{z_{i}\sim S}[\nabla_{\theta}L(z_{i},\theta_{S}^{*})\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}].$ (3)
@@
$\tau_{\mathrm{IFFIM},\lambda,P}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=-\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})^{\top}P(P^{\top}F_{S}P+\lambda I_{\tilde{p}})^{-1}P^{\top}\nabla_{\theta}L(z_{i},\theta_{S}^{*}).$ (4)
Another popular method, TracIn [30], computes the influence of $z_{i}$ on $z^{\prime}$ by tracing the entire optimization process based on stochastic gradient descent. Mathematically, it works as if substituting $H_{S}^{-1}$ with the identity matrix, while taking average across training model checkpoints weighted by the learning rate: $\tau_{\mathrm{TracIn}}(z^{\prime},z_{i})\mathrel{\mathop{\ordinarycolon}}=\sum_{t}\eta_{t}\nabla_{\theta}f(z^{\prime},\theta_{t})^{\top}\nabla_{\theta}L(z_{i},\theta_{t}),$ where $\theta_{t}$ and $\eta_{t}$ are the parameters and learning rate at checkpoint $t$.
@@ 2.2 Evaluation Metrics for Data Attribution
$c_{s}(\tau,z^{\prime},\overline{A})\mathrel{\mathop{\ordinarycolon}}=\mathrm{SpearmanCorr}\bigg(\{f(z^{\prime},\theta^{*}_{A_{j}})\}_{j\in[s]},\big\{\sum_{z\in A_{j}}\tau(z^{\prime},z)\big\}_{j\in[s]}\bigg),$ (5)
@@ 3.1 Experimental Setup
We evaluate a range of data attribution methods, including IF and its variants (CG and LiSSA) [22], TracIn [30], TRAK [29], and LoGra [6], introduced in Section˜2.1.
@@
We perform empirical evaluations on three standard benchmark settings: MNIST [25] with a multilayer perceptron (MLP), CIFAR-2 [24] with ResNet-9 [14], and WikiText2 [28] with the GPT2 [31], which consists of both image and text data with various model sizes. We use the implementation from the dattri library [8] for the datasets and models.
@@
We consider both the common hyperparameters introduced in Section˜2.1 and some critical method-specific hyperparameters. For TRAK, we experiment with regularization, projection-dimension, and training-epoch. For TracIn, we search for projection-dimension, normalization, and checkpoint-selection. For IF, we analyze regularization and training-epoch, as well as max-iteration for the CG variant, and scaling and recursion-depth for the LiSSA variant. For LoGra, we search for regularization, projection-dimension, and training-epoch. We design the search space for each hyperparameter around its default value proposed by the original papers. The detailed definitions and the search space of the hyperparameters are stated in Appendix A.1.
@@ 3.2 Experimental Results
((c)) CIFAR-2+ResNet-9. Attributor: TRAK. HP: projection-dimension, regularization.
@@
((e)) CIFAR-2+ResNet-9 & MNIST+ MLP. Attributor: TRAK. HP: training-epoch.
@@
[…]
@@ 4 A Case Study on the Regularization Term in Influence Function
4.1 Problem: Hyperparameter Selection of $\lambda$ in Influence Function
We study the problem of selecting $\lambda$ using the IFFIM version of influence function, $\tau_{\mathrm{IFFIM},\lambda}$, as defined in Eq.˜2.
@@ 4.1 Problem: Hyperparameter Selection of $\lambda$ in Influence Function
Recall Definition˜2.1, where we have denoted LDS as $c_{s}(\tau,z^{\prime},\overline{A})$ for an attributor $\tau$, a test sample $z^{\prime}$, and samples of subsets $\overline{A}$. The hyperparameter selection problem can be formulated as the following optimization problem:
@@
$c_{p}(\tau,z^{\prime})\mathrel{\mathop{\ordinarycolon}}=\mathrm{PearsonCorr}\bigg(\{f(z^{\prime},\theta^{*}_{A})\}_{A\subset S,|A|=a},\big\{\sum_{z\in A}\tau(z^{\prime},z)\big\}_{A\subset S,|A|=a}\bigg).$ (6)
Note that Spearman correlation is equivalent to the Pearson correlation of the rank-transformed variables. Therefore, both statistics measure monotonic relationships and are generally expected to be highly correlated [10]. The new objective, Population Pearson LDS, is differentiable and replaces the sampled subsets $\overline{A}$ with all size $a$ subsets of $S$, which makes it easier to analyze. Specifically, we can define the partial derivative of $c_{p}$ with respect to $\lambda$ as
$\dot{c_{p}}(\lambda;z^{\prime})\mathrel{\mathop{\ordinarycolon}}=\frac{\partial c_{p}(\tau_{\mathrm{IFFIM},\lambda},z^{\prime})}{\partial\lambda}.$ (7)
@@
In our hyperparameter sensitivity experiments in Section˜3, we observe that, empirically, the LDS with respect to the regularization term $\lambda$ often has a uni-modal and concave shape. This observation turns our attention from the original maximization problem to the analysis of the partial derivative $\dot{c_{p}}(\lambda;z^{\prime})$. We first show in the following Lemma˜4.2 that our empirical observation about the shape of the LDS as a function of $\lambda$ is not a coincidence.
@@
Assuming the test sample $z^{\prime}$ follows a similar distribution as the training data, under some technical assumptions deferred to Section˜B.4, [^7] This involves non-trivial technical assumptions that do not always strictly hold. But we have made intuitive and empirical justifications about them. with high probability over the choice of test sample $z^{\prime}$, there exists some $C>0$ such that for $0<\lambda<C$,
@@
Lemma˜4.2 suggests that for a range of small $\lambda$ starting from 0, $c_{p}$ as a function of $\lambda$ is increasing. With the observation regarding the shape of the LDS curve, ideally, we can solve for the optimal hyperparameter $\lambda^{*}$ as the stationary point with $\dot{c_{p}}(\lambda^{*};z^{\prime})=0$. However, the partial derivative $\dot{c_{p}}(\lambda^{*};z^{\prime})$ still involves quantities $\theta_{A}^{*}$ for $A\subset S$, which require model retraining.
4.2 Analysis: A Sufficient Condition for Monotonicity in $\lambda$
Instead of directly solving for the stationary point, we derive a sufficient condition (Theorem˜4.3) for the derivative to be positive, which leads to a surrogate indicator that does not require model retraining (see Definition˜4.6 in Section˜4.3). Before we state this condition, we first introduce some key notations and their intuitive interpretations.
LOO weighted loss gradients.
We let $D_{a}$ denote the uniform distribution of size $a$ subsets of $S$. Define $g_{z^{\prime}}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{n}\sum_{i=1}^{n}\alpha_{z^{\prime},i}\nabla_{\theta}L(z_{i},\theta_{S}^{*})\in\mathbb{R}^{p}$, where $\alpha_{z^{\prime}}\in\mathbb{R}^{n}$ with $\alpha_{z^{\prime},i}\mathrel{\mathop{\ordinarycolon}}=\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})|z_{i}\in A]-\mathbb{E}_{A\sim D_{a}}[f(z^{\prime},\theta_{A}^{*})]$. As the name suggests, we point out that $\alpha_{z^{\prime},i}$ can be interpreted as the LOO influence of $z_{i}$ on $z^{\prime}$ under certain conditions. Under this interpretation, $g_{z^{\prime}}$ becomes the sum of loss gradients of training points weighted by their LOO influences on the test example $z^{\prime}$. More details can be found in Section˜B.2.
Other necessary notation.
For simplicity of notation in our subsequent analysis, we let $v_{z^{\prime}}$ denote the gradient $\nabla_{\theta}f(z^{\prime},\theta_{S}^{*})$, and define several bilinear terms of $v_{z^{\prime}}$, $g_{z^{\prime}}$, or $\alpha_{z^{\prime}}$. They will be used to measure and compare different $\theta$ update directions (see Theorem˜4.3 and its interpretations below). Denote $t_{k,z^{\prime},\lambda}\mathrel{\mathop{\ordinarycolon}}=v_{z^{\prime}}^{\top}(F_{S}+\lambda I_{p})^{-k}F_{S}v_{z^{\prime}}$. Further, we let $r_{z^{\prime},\lambda}\mathrel{\mathop{\ordinarycolon}}=-v_{z^{\prime}}^{\top}(F_{S}+\lambda I_{p})^{-1}g_{z^{\prime}}$ and $o_{z^{\prime},\lambda}\mathrel{\mathop{\ordinarycolon}}=\alpha_{z^{\prime}}^{\top}(JJ^{\top}+n\lambda I_{n})^{-1}\alpha_{z^{\prime}}$, where $J\in\mathbb{R}^{n\times p}$ has its $i$th row being $\nabla_{\theta}L(z_{i},\theta_{S}^{*})^{\top}$. The relationship between $(F_{S}+\lambda I_{p})^{-1}$ and $(JJ^{\top}+n\lambda I_{n})^{-1}$ is embodied in Lemma˜B.3. Intuitively, these terms define semi-inner products of $v_{z^{\prime}}$, $(F_{S}+\lambda I_{p})^{-1}v_{z^{\prime}}$, $g_{z^{\prime}}$ and $\alpha_{z^{\prime}}$ weighted by matrices related to the FIM.
Now we are ready to state a sufficient condition for the monotonicity of Population Pearson LDS.